Joe has a collection of nickels and dimes that is worth $5.65. If the number of dimes are doubled and the number of nickels were increased by 8, the value of the coins would be $10.45. How many dimes does he have?
Answer by stanbon(75887)

Joe has a collection of nickels and dimes that is worth $5.65. If the number of dimes are doubled and the number of nickels were increased by 8, the value of the coins would be $10.45. How many dimes does he have?
--------------------
Value equation: 5n + 10d = 565 cents
value equation:5(n+8)+10(2d)=1045 cents
-------------------------------------------
Rearrange:
5n + 10d = 565
5n + 20d = 1005
------------------------
10d = 440
d = 44 (# of dimes)
----
Substitute to solve for "n":
5n + 440 = 565
5n = 125
n =25 (# of nickels)
========================
Cheers,
Stan H.